OpenStudy (anonymous):

what is the derivative of f(x)=ln((x^(8)-8)^(5))

OpenStudy (anonymous):

The derivative of the natural log function is \[\frac{d}{dx}[\ln u(x)]=\frac{u'(x)}{u(x)}\] In this case, \(u(x)=\left(x^8-8\right)^5\). However, you can use a log property to rewrite the function: \[f(x)=\ln\left(x^8-8\right)^5\\ f(x)=5\ln\left(x^8-8\right)\] So, you have \[f'(x)=5\frac{\frac{d}{dx}[x^8-8]}{x^8-8}\]

OpenStudy (anonymous):

thank you so much!!

OpenStudy (anonymous):

You're welcome